solve 5/x + x/x+2 = 4/x^2+2x
Answer by checkley77(12569)

5/x + x/(x+2) = 4/(x^2+2x)
5/x+x/(x+2)=4/x(x+2) combine the left side terms.
[5(x+2)+x^2]/x(x+2)=4/x(x+2) multiply both sides by x(x+2)
5(x+2)+x^2=4
5x+10+x^2=4
x^2+5x+10-4=0
x^2+5x+6=0
(x+3)(x+2)=0
x+3=0
x=-3 ans.
x+2=0
x=-2 ans.